On the first part of a road trip to Batangas travelling on the highway, Kelly averages 60 kph. On the rest of the roadtrip, which was 10 kilometers longer than the first part, she averaged 50 kph. Find the total distance to Batangas if the second part of the trip took 30 minutes more than the first.
Help please!

Answer by rothauserc(4718)

We are given two formulas for rate * time = distance
60*t = d1, where t is time and d1 is distance for first part of trip
50 *(t+(1/2)) = d1+10, where (1/2) is half an hour or 30 minutes
substitute for d1 in second equation
50 *(t+(1/2)) = (60*t)+10
50t +25 = 60t +10
10t = 15
t = 15/10 = 1.5 hours, therefore
60*1.5 = 90 kilometers and
90 + 10 = 100 kilometers and
90 + 100 = 190 kilometers
now 190 kilometers is the total distance